Ticket Number: HEA-NOC/20060427-3 Ticket Status: CLOSE
Ticket Type: unscheduled Resolver: HEA-NOC
Ticket Opened: 20060427 09:57 UTC+1 Problem Start: 20060427 06:30 UTC+1
Ticket Closed: 20060502 10:41 UTC+1 Problem End: 20060427 09:30 UTC+1
Site/line: HEA-NOC, mail
Mail services to and from the HEAnet NOC were interupted for roughly one
hour and were suffering intermittent issues from around 06:30 UTC +1.
Attempted SMTP connections were temporarily rejected. All mail should
resend depending on individual MTA configurations. If you are concerned
about mail sent this morning (prior to 09:40 UTC +1) then please follow-up
with the NOC, however there should be no need to resend.
20060427 09:45 BN Determined problem with hanging processes, cleaned up and
confirmed that mail was flowing in and out of mail.heanet.ie once more.
20060428 08:55 SD Mail services affected again this morning, same as the
previous morning, had to restart sohpie (AV scanner) to get it going
20060502 10:40 BN Problem has not recurred and while investigations are
ongoing this public ticket is being closed.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20060427-3